Lewis Auto Company manufactures a part for use in its production of automobiles. When 10,000 units are produced, the costs per unit are:

Direct materials $15

Direct manufacturing labor 60

Variable manufacturing overhead 26

Fixed manufacturing overhead 32

Total $133

Monty Company has offered to sell to Lewis Auto Company 10,000 units of the part for $122 per unit. The plant facilities could be used to manufacture another item at a savings of $182,000 if Lewis accepts the offer. In addition, $24 per unit of fixed manufacturing overhead on the original part would be eliminated.

Required: Which alternative is best for Lewis Auto Company? By how much – please show your work!
Answer #1

Relevant cost = Direct Material + direct manufacturing + Variable manufacturing overhead + Relevant fixed cost
= 15+60+26+20

= 122

Cost if the part is bought
Purchase price 10,000*122 = 1,220,000

Add: Unavoidable fixed cost 10,000*32 = 320,000 = 1,540,000